Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.) μ = 14.0; σ = 4.2 P(10 ≤ x ≤ 26)
Answer #1

Solution:

Given that,

mean =\mu  = 14.0

standard deviation =\sigma = 4.2

p ( 10 \leq x \leq 26 )

= p ( 10- 14.0 / 4 .2)\leq ( x -  \mu/ \sigma) \leq ( 26 - 14.0 / 4.2 )

= p ( - 4 / 4.2 \leq z \leq 12 / 4.2)

= p (- 0.95 \leq z \leq 2.86 )

= p (z \leq 2.86 ) - p ( z \leq - 0.95 )

Using z table

= 0.9979 - 0.1711

= 0.8268

Probability = 0.8268